OpenStudy (anonymous):

Friends , plse explain its answer a point on the curve 2y^3 + x^2 = 12y at which the tangent to the curve is vertical is

OpenStudy (anonymous):

vertical line has a slope of ....?

OpenStudy (anonymous):

infinity

OpenStudy (anonymous):

good. now, tangent slope is obtained by....?

OpenStudy (anonymous):

dy/dx

OpenStudy (anonymous):

perfect. so, lets find it

OpenStudy (anonymous):

implicit diff

OpenStudy (anonymous):

dy/dx=x/(6-3y^2)

OpenStudy (anonymous):

yep.. now, when does this slope become infinite?

OpenStudy (anonymous):

hint: vertical asymptote

OpenStudy (anonymous):

@msingh well?

OpenStudy (anonymous):

i have done wrong

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

is this answer

OpenStudy (anonymous):

\[ 6-3y^2=0\implies y=\pm\sqrt{2} \] for those value "y", you have a tangent to the graph that is vertical. -> two of them

OpenStudy (anonymous):

we have to put coeff of y = 0 when tangent is vertical

OpenStudy (anonymous):

how? and where?

OpenStudy (anonymous):

when we have calculate dy/dx, what after that

OpenStudy (anonymous):

ok, then you have x^2=0...

OpenStudy (anonymous):

but, explain your statement geometrically

OpenStudy (anonymous):

x or x^2=0

OpenStudy (anonymous):

so, what is your answer.?

OpenStudy (anonymous):

nahi pata

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

Then follow the steps.. math does not lie

OpenStudy (anonymous):

k

OpenStudy (anonymous):

dy/dx=x/(6-3y^2) then 6-3y^2=0 y=+-under root 2 because, for those value "y", you have a tangent to the graph that is vertical. -> two of them

OpenStudy (anonymous):

till here ,is it correct

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

so, you know the ordinates at which the tangent goes vertical. now use f(x,y) and find "x" values those two (may be 4, but some outside the domain) points will give you the required points

OpenStudy (anonymous):

x^2=4 underrot 2, when y=+underroot 2

OpenStudy (anonymous):

x^2=-4 underrot 2, when y=-underroot 2

OpenStudy (anonymous):

no check again.. 8sqrt(2)

OpenStudy (anonymous):

we have to put the value of y in the given equation 2y^3+x^2=12y

OpenStudy (anonymous):

yup u r right 8 underroot 2

OpenStudy (anonymous):

right.. so, now, take sq-root, you'd get one +ve and one -ve, that is your answer.. the two (x,y) points

OpenStudy (anonymous):

(2under root 2 under root 2,under root 2)

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

iguess.. one +ve x and one -ve "x" should be symmetric/

OpenStudy (amistre64):

infinite slope? oy vey!!

OpenStudy (amistre64):

assuming that y is a function of x ... implicitly. that values of x that produce a tangent that is "undefined" will indicate vertical slopes

OpenStudy (amistre64):

2y^3 + x^2 = 12y 6y^2 y' + 2x x' = 12; dx/dx=x' = 1 6y^2 y' + 2x = 12 6y^2 y' = 12 - 2x y' = (12 - 2x)/6y^2

OpenStudy (amistre64):

opps, i made an error :) 2y^3 + x^2 = 12y 6y^2 y' + 2x x' = 12 y'; dx/dx=x' = 1 ^^ forgot to pop out a y' :) 2x = 12 y' - 6y^2 y' x = 6 y' - 3y^2 y' x = (6 - 3y^2) y' x/(6-3y^2) = y' thats better

OpenStudy (amistre64):

so, your derivative is correct ... thats good so this y' becomes undefined when we divide by 0 6-3y^2 = 0, when y=??

OpenStudy (amistre64):

is say when y=+- sqrt(2)

OpenStudy (amistre64):

so far youve done good, then you plugged that into the original setup to find teh x values

OpenStudy (amistre64):

2y^3 + x^2 = 12y 2sqrt(2)^3 + x^2 = 12sqrt(2) x^2 = 12sqrt(2) - 4sqrt(2) x^2 = 8sqrt(2) x = +- sqrt(8sqrt(2)) x = +- 2sqrt(2sqrt(2)) looks good to me

OpenStudy (anonymous):

@amistre64 thankk you

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

@amistre64 so, what is the difference? about using the words "infinite" vs "undefined"?

OpenStudy (amistre64):

just semantics :) one method ignores mathing principles, the other applies them. \[infinite ~slope:\frac{\infty}{1}\]\[undefined~slope:~\frac{k}{0}\]

OpenStudy (amistre64):

of course you could have undefined slopes that use other undefinable function :) ln(0)/3 is undefined along the reals 2/(2n)rt(-6) is undefined along the reals etc.....

OpenStudy (anonymous):

@even semantically, you are infinitesimally closer to "y" by not there, technically, you the ratio is still defined. Also, in Algebra, you say it is undefined. In geometry, it is "usually" safe to take it to define infinity.

OpenStudy (anonymous):

Reference-> "1,2,3.. Infinity by George Gamow"

OpenStudy (amistre64):

Gamow is a hack!! lol

OpenStudy (anonymous):

@amistre64 lol. and who doesnt wish to be a membor of the whacky group!

OpenStudy (amistre64):

... bob dole, or so ive been led to spread that rumor about

OpenStudy (anonymous):

@amistre64 lmao

OpenStudy (anonymous):

@amistre64 Drink to that brother.